# Emotion Recognition Using Transformers with Masked Learning

### 19 Mar 2024 v1

@@ 1 Introduction
Recently, deep learning has undergone significant changes in various fields such as computer vision, natural language processing, and especially in analyzing human emotions and behaviors. One of the key developments in this field is the Affective Behavior Analysis in-the-wild (ABAW) competition held by Kollias et al. [14, 9, 13, 12, 10, 11, 15, 6, 7, 16, 8, 22] These competitions facilitate research by providing diverse and challenging datasets such as AffWild2, C-EXPR-DB, and Hume-Vidmimic2, encouraging the development of models capable of accurately assessing complex emotional states. These models provide keys through Valence-Arousal (VA) estimation, facial expression recognition, and Action Unit (AU) detection, which are essential components in understanding human emotions.

### 23 Mar 2024 v2

@@ 1 Introduction
Recently, deep learning has undergone significant changes in various fields such as computer vision, natural language processing, and especially in analyzing human emotions and behaviors. One of the key developments in this field is the Affective Behavior Analysis in-the-wild (ABAW) competition held by Kollias et al. [14, 9, 13, 12, 10, 11, 15, 6, 7, 16, 8, 23, 17] These competitions facilitate research by providing diverse and challenging datasets such as AffWild2, C-EXPR-DB, and Hume-Vidmimic2, encouraging the development of models capable of accurately assessing complex emotional states. These models provide keys through Valence-Arousal (VA) estimation, facial expression recognition, and Action Unit (AU) detection, which are essential components in understanding human emotions.